OpenStudy (anonymous):

Find the inverse of the function: f(x)=(5)/(1-3x) I am not really sure how to do this I have done y=(5)/(1-3x) and attempted to make x the subject but have had no success. Thanks for any help!

OpenStudy (azureilai):

When you try to find the inverse, you switch the place of the x and y in a function, so it would be x=5/(1-3y) now isolate y to one side of the equation

OpenStudy (anonymous):

Yes I know that but I'm not so sure how to isolate y... Would it be something like 5/x=1-3y?

OpenStudy (azureilai):

x(1-3y)=5 distribute x x-3xy=5 simplifu (5-x)=-3xy (5-x)/(-3x)=y

OpenStudy (anonymous):

The answers say its (-5)/((1-3x)^2) x -3 Is this the same as your answer?

OpenStudy (alexwee123):

f(x)=(5)/(1-3x) F^-1(x)=> x=5/(1-3y) x(1-3y)=5 1-3y=5/x -3y=5/x -1 y=-(5/3x)+1/3 y=(-5+x)/3x y=(5-x)/(-3x) @Azureilai is right

OpenStudy (anonymous):

Ok thank you